OpenStudy (anonymous):

solve square root of 3x-7=square root of 2x+5

OpenStudy (rock_mit182):

hi :\[\sqrt{3x-7}=\sqrt{2x+5}\]

OpenStudy (rock_mit182):

First: square both sides of the equation

OpenStudy (anonymous):

what is the answer. my numbers don't equal eachother, I squared both sides I got x=74 but that is wrong

OpenStudy (rock_mit182):

let me see

OpenStudy (rock_mit182):

that's cause i guess you made a mistake in solving the eqtuation. let me see your work

OpenStudy (rock_mit182):

so we'LL be able to find the error

OpenStudy (anonymous):

I figured it out, I squared the number twice, the answer is 12 correct?

OpenStudy (rock_mit182):

that's correct !

OpenStudy (rock_mit182):

;)

OpenStudy (anonymous):

thanks for your help,.